HUBBLE SPACE TELESCOPE DAILY REPORT #1200 PERIOD COVERED: 0000Z Daily Status Report as of 223/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOC/96 5526 (FOC UV Throughput Monitoring: Cycle 4) The Faint Object Camera (f/96) was used to make calibration observations of ultraviolet standard star BPM16274, to monitor the ultraviolet throughput of the HST-FOC. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3: Darks) The WF/PC-2 was used to obtain several 30-minute long, dark exposures to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed FGS 5556 (Binary Plate Scale: Cycle 4 Calibration) The Fine Guidance Sensors were used to make observations of standard star HD16620 to study the plate scale calibration in the central portion of FGS #3. This will be used in binary star observations. The observations were completed as planned, and no problems were reported. 1.4 Completed WF/PC-2 5091 (GTO High Latitude Parallel Program: Cycle 4) The WF/PC-2 was used in parallel with the prime science observations to make high galactic latitude observations greater than +/-30 degrees. The observations were completed as scheduled, and no problems were reported 1.5 Completed WF/PC-2 5143 (Observations of Galaxies Underlying Quasars) The WF/PC-2 was used to make observations of the low redshift Quasar PKS2128-12 to study the morphology of the underlying galaxy. The observations were executed as scheduled, and no problems were reported. 1.6 Completed WF/PC-2 5571 (WF/PC-2 Cycle 4 Calibration: Earthflats, Smaller Set) The WF/PC-2 was used to make several ultraviolet and narrow-band filter Earth flat calibration observations to correct for the telescope's illumination pattern. The observations were completed as planned, and no problems were reported. 1.7 Completed GHRS/1 5642 (Observations of the Impact of Comet Shoemaker-Levy (1993E) and Jupiter) The High Resolution Spectrograph was used to make made observations of the planet Jupiter to support the Jupiter/ Shoemaker-Levy impact event. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled :12 Successful :12 2.3 Significant HSTARs: None 2.4 Operations Notes: No problems were reported on any of the spacecraft subsystems during the last 24 hours. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: To minimize the possiblity of damage from the Perseid metior stream, the telescope will be pointed near the anti-radiant point starting at 224/0330Z until 224/1420Z. In addition, to minimize the solar array cross section to the stream, the solar arrays will be moved from their normal position. /DMH -END-